Both you and your older brother would like to have $22,000 in 14 in years. Because of your success in this class, you feel that you are a more savvy investor than your brother and will be able to earn an annual return of 10.8 percent compared to your brother's 9.3 percent. How much less than your brother will you have to deposit today?
Answer #1

Present value=cash flow*Present value of discounting factor(rate%,time period)

For me:

Present value=22000/1.108^14

=22000*0.237926795

=$5234.39

For my brother:

Present value=22000/1.093^14

=22000*0.28795043

=$6334.91

Hence less deposit =$6334.91-$5234.39

=$1100.52(Approx).